# Dynamical black hole entropy beyond general relativity from the Einstein frame

### v2

@@ 3 Dynamical black hole entropy in general relativity
where $M$ and $J$ are ADM definitions of the mass and angular momentum of the spacetime, respectively. This is because $\Delta\delta S_{\text{dyn}}=0$ for source-free perturbations $\delta T_{ab}=0$, $S_{\text{dyn}}$ equals the usual Bekenstein-Hawking entropy at the bifurcation surface, and it has been proven that Bekenstein-Hawking entropy satisfies the comparison first law at the bifurcation surface. It is noteworthy that this relation still holds if matter fields are present as long as the combined Einstein-matter system admits a Hamiltonian formulationWald:1995yp ; Sudarsky:1992ty .
In the case of vacuum perturbations there are no external matter fields and $\delta T_{ab}=0$. Then (22) indicates that the dynamical black hole entropy is a constant at first order in perturbation theory. So we have to keep track of the second order in perturbation theory to study the leading-order behavior of the dynamical black hole entropy. Hollands, Wald and Zhang obtained the leading-order behavior in that case based on the varied fundamental identity of the covariant phase space formalismHollands:2024vbe . In what follows we are going to present a different derivation of the second-order behavior of $S_{\text{dyn}}$ by considering the second variation of the Raychaudhuri equationRignon-Bret:2023fjq .
@@
Thus the perturbed expansion $\delta\theta$ is a constant along the null generators of $\mathcal{H}^{+}$. Recalling that $\theta$ vanishes at future infinity in the perturbed spacetime due to the teleological definition of the event horizon, the perturbed expansion $\delta\theta$ must vanish on $\mathcal{H}^{+}$ollands:2012sf . We then consider the second-order variation of (14) to obtain
@@
We multiply this equation on both sides by $\kappa v$, integrate over the future horizon between $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$, and integrate the left hand side by parts in the same way shown above
$\kappa\int_{\mathcal{C}(v_{1})}\mathrm{d}A\left[v\delta^{2}\theta\right]_{v_{1}}^{v_{2}}-\kappa\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \delta^{2}\theta=-2\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa v\left(\delta\sigma_{ab}\delta\sigma^{ab}\right).$ (32)
What’s more, on the left hand side of this equation we may pull $\delta^{2}$ to the front of the integrals as both $\theta$ and $\delta\theta$ vanish. Finally the result is
@@
Our result agrees with that derived by varying the fundamental identity of the covariant phase space formalismHollands:2024vbe . The term $\delta\sigma_{ab}\delta\sigma^{ab}$ can be interpreted as the energy flux of the weak gravitational wavesAshtekar:2021kqj . Since $\delta\sigma_{ab}\delta\sigma^{ab}\geq 0$, the second law holds at second order for vacuum perturbations of GR.
4 Dynamical black hole entropy in $f(R)$ gravity
In this section we derive both the non-stationary physical process first law and the non-stationary comparison first law for $f(R)$ gravity using the Einstein frame. Our results and and the dynamical black hole entropy formula for $f(R)$ gravity agree with those derived by means of the Noether charge methodHollands:2024vbe ; Visser:2024pwz . For vacuum perturbations we also study the second-order behavior of the dynamical black hole entropy, and observe that the second law is obeyed both in the Einstein frame and in the original frame. What’s more, we determine the relationship between the dynamical black hole entropy of the cross-section $\mathcal{C}$ and the Wald entropy of the generalized apparent horizon $\mathcal{T}$ in the original frame.
4.1 The conformal transformation and the Einstein frame
@@
And in this paper, we assume that [^3] We impose $f^{\prime}(R)>0$ to ensure that the effective gravitational coupling strength $G_{\text{eff}}=G/f^{\prime}(R)$ for the $f(R)$ theory is positiveSotiriou:2008rp ; Nojiri:2017ncd . On the other hand, if $f^{\prime\prime}(R)=0$, then $f(R)$ is linear to $R$, and the resulting theory is reduced to GR.
@@
because with the on-shell condition $\varphi=R$, this action reduces to the original one. To utilize the conclusions of the last section, we turn to the Einstein frame and redefine the scalar field as followsBarrow:1988xh ; Whitt:1984pd
$g^{E}_{ab}\equiv\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}g_{ab},\quad\phi\equiv\frac{1}{\sqrt{16\pi G_{D}}}\sqrt{\frac{2\left(D-1\right)}{D-2}}\ln f^{\prime}(\varphi).$ (37)
@@
Since the conformal transformation preserves the causal structure of the spacetime, and $g^{E}_{ab}$ is asymptotically flat as long as $g_{ab}$ is asymptotically flat, the conformal transformation (37) of an asymptotically flat black hole spacetime remains an asymptotically flat black hole spacetime with the same event horizon, at least the conformal factor is regular on the horizonJacobson:1993pf ; Jacobson:1995uq . Moreover, the conformal factor $\omega=\left[f^{\prime}(\varphi)\right]^{\frac{1}{D-2}}=\left[f^{\prime}(R)\right]^{\frac{1}{D-2}}$ is independent of the Killing time $\tau$ in the stationary black hole spacetime, so the conformal transformation (37) of a stationary black hole solution is still a stationary black hole solution, with the horizon Killing field given by $\xi^{a}$ up to a normalization factor. If the horizon Killing field $\xi^{a}$ is normalized as $\xi^{a}=(\partial_{t})^{a}+\Omega_{\mathcal{H}}(\partial_{\theta})^{a}$ with $(\partial_{t})^{a}$ representing time translations at infinity in the original frame, we normalize $\xi_{E}^{a}=\alpha\xi^{a}$ with the normalization constant $\alpha$ determined by
@@
Therefore the relation between $\kappa^{E}$ and $\kappa$ reads $\kappa^{E}=\alpha\kappa=\left[f^{\prime}(0)\right]^{-\frac{1}{D-2}}\kappa$. After the conformal transformation, the null affine parameter $v^{E}$ on the horizon in the Einstein frame is related to that in the original frame $v$ byWald:1984rg
$\frac{\mathrm{d}v^{E}}{\mathrm{d}v}=c\left[f_{R}(v,x)\right]^{\frac{2}{D-2}},$ (42)
where $c$ is an arbitrary constant and $f_{R}(v,x)$ is the value of $f^{\prime}(\varphi)=f^{\prime}(R)$ at $(v,x)$ on the horizon. As a result,
$v^{E}=c\int_{0}^{v}\mathrm{d}v^{\prime}\left[f_{R}(v^{\prime},x)\right]^{\frac{2}{D-2}}.$ (43)
Since $0=\partial_{\tau}f^{\prime}(R)=\kappa v\partial_{v}f_{R}(v,x)$ in the stationary background, $f_{R}(v,x)$ does not depend on $v$, so $v^{E}$ is proportional to $v$ on the horizon of the stationary black hole.
4.2 The physical process first law
Since null geodesics are conformally invariant and the Einstein equation holds for $g^{E}_{ab}$ in the Einstein frame, the derivation of Sec.3 is still tenable for $f(R)$ gravity in the Einstein frame. Let
@@ 4.2 The physical process first law
denote the sum of the stress-energy tensor of the external matter fields in the Einstein frame $T^{E}_{ab}=-\frac{2}{\sqrt{-g^{E}}}\frac{\delta S_{M}}{\delta g_{E}^{ab}}=T_{ab}/f_{R}(v,x)$ as well as that of the auxiliary scalar field. Notice that
@@
where we have used the fact that $k_{E}^{a}\partial_{a}\phi=0$ and $\left.g^{E}_{ab}k^{a}k^{b}\right|_{\mathcal{H}^{+}}=0$ in the stationary background as well as the gauge condition $\delta g_{ab}k_{E}^{a}k_{E}^{b}\overset{\mathcal{H}^{+}}{=}0$ for the perturbations. Thus, only the variation of the external matter fields contributes to the first-order variation of $\mathcal{T}_{ab}$. For perturbations satisfying the gauge conditions listed in Sec.2 that fix $k_{E}^{a}$ on the horizon, the non-stationary physical process first law for $f(R)$ gravity in the Einstein frame reads
@@
Next we would like to recast the physical process first law and the formula for the dynamical black hole entropy in the original frame. As the perturbation performed on the black hole is non-stationary, the gauge conditions that fix $k_{E}^{a}$ will not generally fix $k^{a}$ in the original frame. Since for non-stationary perturbations, $f_{R}(x,v)$ is not necessary a constant along the null geodesics on $\mathcal{H}^{+}$ and $v^{E}$ presumably would not be proportional to $v$ accordingly. While as mentioned in Section 2, the essence of those gauge conditions lies in how to identify spacetime points in two slightly different spacetimes. Those gauge conditions that fix $k_{E}^{a}$ under the perturbations imply that we correspond points with the same $v^{E}$ in the two spacetimes. To rewrite (46) and (47) in the original frame, in what follows we are going to compare spacetime points sharing the same $v$ on the horizon, and replace $v^{E}$ with $v$ in the formula of $S^{E}_{\text{dyn}}$.
@@
Without loss of generality, in the following discussion we set $v^{E}=v$ in the stationary background, and perform the conformal transformation (37) on the perturbed geometry $g_{ab}+\delta g_{ab}$ to obtain the non-stationary black hole solution in the Einstein frame. Then $v^{E}=v+\delta V(x,v)$ on the horizon according to (43) in the perturbed geometry, where $\delta V(x,v)$ is a correction function derived from expanding the conformal factor $\omega^{2}=\left[f_{R}(x,v)\right]^{\frac{2}{D-2}}$ with respect to the non-stationary perturbation of the metric, and it is of first order in perturbation theory. To begin with, let $F(v,x)$ represent a quantity on the horizon, such as the entropy $S_{\text{dyn}}$ or the component of the stress-energy tensor $T_{vv}(v,x)$, which is stationary in the background, and let $\tilde{F}(v,x)$ denote the corresponding quantity in the perturbed spacetime, then $\delta F(v,x)=\tilde{F}(v,x)-F(v,x)$ does not change to first order as we switch from $v^{E}-$identification to $v-$identification, because
$\delta_{v^{E}}F-\delta_{v}F=\tilde{F}(v+\delta V,x)-\tilde{F}(v,x)\sim\partial_{v}\tilde{F}\ \delta V=\mathcal{O}(\epsilon^{2}),$ (48)
@@
where $\delta_{v}$ stands for the comparison between points with the same $v$ in the stationary background and in the perturbed spacetime, and the same connotation holds for $\delta_{v^{E}}$. We have also used that $\partial_{v}\tilde{F}$ is of first order in the perturbed spacetime as $F$ is stationary in the unperturbed background. Therefore (46) does not change to first order under the re-identification of spacetime points. In addition, although the difference $\delta V$ between $v$ and $v^{E}$ leads to first-order corrections to the integral over $v$ and to the null generator $k^{a}$ in the perturbed geometry, the right hand side of (46) will not change to first order if we replace $k^{a}$ by $k_{E}^{a}$ and replace $\int_{v^{E}_{1}}^{v^{E}_{2}}\mathrm{d}v^{E}$ by $\int_{v_{1}}^{v_{2}}\mathrm{d}v$, since $\delta T^{E}_{ab}$ is a first-order small quantity and this substitution leads to second-order corrections. Moreover, in the perturbed geometry,
@@
$\begin{split}S^{E}_{\text{dyn}}&=\frac{1}{4G_{D}}\left(1-v^{E}\frac{\mathrm{d}v}{\mathrm{d}v^{E}}\frac{\mathrm{d}}{\mathrm{d}v}\right)\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}}\\ &=\frac{1}{4G_{D}}\left[1-\left(v+\delta V\right)\left(1-\frac{\mathrm{d}\delta V}{\mathrm{d}v^{E}}\right)\frac{\mathrm{d}}{\mathrm{d}v}\right]\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}}\\ &=\frac{1}{4G_{D}}\left(1-v\frac{\mathrm{d}}{\mathrm{d}v}\right)\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}}+\mathcal{O}(\epsilon^{2}),\end{split}$ (49)
@@
Our results agree with those obtained by means of the Noether charge methodHollands:2024vbe ; Visser:2024pwz . Moreover, if the stress-energy tensor of the external matter field satisfies the null energy condition $\delta T_{ab}k^{a}k^{b}\geq 0$, the dynamical black hole entropy will obey the linearized second law $\Delta\delta S_{\text{dyn}}\geq 0$ under the perturbation.
4.3 The comparison first law

### v4

@@ III Dynamical black hole entropy in general relativity
where $M$ and $J$ are ADM definitions of the mass and angular momentum of the spacetime, respectively. This is because $\Delta\delta S_{\text{dyn}}=0$ for source-free perturbations $\delta T_{ab}=0$, $S_{\text{dyn}}$ equals the usual Bekenstein-Hawking entropy at the bifurcation surface, and it has been proven that Bekenstein-Hawking entropy satisfies the comparison first law at the bifurcation surface. It is noteworthy that this relation still holds if matter fields are present as long as the combined Einstein-matter system admits a Hamiltonian formulation[16, 17].
In the case of vacuum perturbations there are no external matter fields and $\delta T_{ab}=0$. Then (22) indicates that the dynamical black hole entropy is a constant at first order in perturbation theory. So we have to keep track of the second order in perturbation theory to study the leading-order behavior of the dynamical black hole entropy. Hollands, Wald and Zhang obtained the leading-order behavior in that case based on the varied fundamental identity of the covariant phase space formalism[1]. In what follows we are going to present a different derivation of the second-order behavior of $S_{\text{dyn}}$ by considering the second variation of the Raychaudhuri equation[13].
@@
Thus the perturbed expansion $\delta\theta$ is a constant along the null generators of $\mathcal{H}^{+}$. Recalling that $\theta$ vanishes at future infinity in the perturbed spacetime due to the teleological definition of the event horizon, the perturbed expansion $\delta\theta$ must vanish on $\mathcal{H}^{+}$[18]. We then consider the second-order variation of (14) to obtain
@@
We multiply this equation on both sides by $\kappa v$, integrate over the future horizon between $\mathcal{C}(v_{1})$ and $\mathcal{C}(v_{2})$, and integrate the left-hand side by parts in the same way shown above
$\kappa\int_{\mathcal{C}(v_{1})}\mathrm{d}A\left[v\delta^{2}\theta\right]_{v_{1}}^{v_{2}}-\kappa\int_{\mathcal{C}(v_{1})}\mathrm{d}A\int_{v_{1}}^{v_{2}}\mathrm{d}v\ \delta^{2}\theta-2\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \kappa v\left(\delta\sigma_{ab}\delta\sigma^{ab}\right).$ (32)
What’s more, on the left-hand side of this equation we may pull $\delta^{2}$ to the front of the integrals as both $\theta$ and $\delta\theta$ vanish. Finally the result is
@@
Our result agrees with that derived by varying the fundamental identity of the covariant phase space formalism[1]. The term $\delta\sigma_{ab}\delta\sigma^{ab}$ can be interpreted as the energy flux of the weak gravitational waves[19]. Since $\delta\sigma_{ab}\delta\sigma^{ab}\geq 0$, the second law holds at second order for vacuum perturbations of GR.
IV Dynamical black hole entropy in $f(R)$ gravity
In this section we derive both the non-stationary physical process first law and the non-stationary comparison first law for $f(R)$ gravity using the Einstein frame. Our results and the dynamical black hole entropy formula for $f(R)$ gravity agree with those derived by means of the Noether charge method. For vacuum perturbations we also study the second-order behavior of the dynamical black hole entropy, and observe that the second law is obeyed both in the Einstein frame and in the original frame. What’s more, we determine the relationship between the dynamical black hole entropy of the cross-section $\mathcal{C}$ and the Wald entropy of the generalized apparent horizon $\mathcal{T}$ in the original frame.
IV.1 The conformal transformation and the Einstein frame
@@
And in this paper, we assume that [^3] We impose $f^{\prime}(R)>0$ to ensure that the effective gravitational coupling strength $G_{\text{eff}}=G/f^{\prime}(R)$ for the $f(R)$ theory is positive[20, 21]. On the other hand, if $f^{\prime\prime}(R)=0$, then $f(R)$ is linear to $R$, and the resulting theory is reduced to GR.
@@
because with the on-shell condition $\varphi=R$, this action reduces to the original one. To utilize the conclusions of the last section, we turn to the Einstein frame and redefine the scalar field as follows[22, 23]
$\begin{split}g^{E}_{ab}\equiv\left[f^{\prime}(\varphi)\right]^{\frac{2}{D-2}}g_{ab},\quad\phi\equiv\frac{1}{\sqrt{16\pi G_{D}}}\sqrt{\frac{2\left(D-1\right)}{D-2}}\ln f^{\prime}(\varphi).\end{split}$ (37)
@@
Since the conformal transformation preserves the causal structure of the spacetime, and $g^{E}_{ab}$ is asymptotically flat as long as $g_{ab}$ is asymptotically flat, the conformal transformation (37) of an asymptotically flat black hole spacetime remains an asymptotically flat black hole spacetime with the same event horizon, at least the conformal factor is regular on the horizon[24, 25]. Moreover, the conformal factor $\omega=\left[f^{\prime}(\varphi)\right]^{\frac{1}{D-2}}=\left[f^{\prime}(R)\right]^{\frac{1}{D-2}}$ is independent of the Killing time $\tau$ in the stationary black hole spacetime, so the conformal transformation (37) of a stationary black hole solution is still a stationary black hole solution, with the horizon Killing field given by $\xi^{a}$ up to a normalization factor. If the horizon Killing field $\xi^{a}$ is normalized as $\xi^{a}=(\partial_{t})^{a}+\Omega_{\mathcal{H}}(\partial_{\theta})^{a}$ with $(\partial_{t})^{a}$ representing time translations at infinity in the original frame, we normalize $\xi_{E}^{a}=\alpha\xi^{a}$ with the normalization constant $\alpha$ determined by
@@
Therefore the relation between $\kappa^{E}$ and $\kappa$ reads $\kappa^{E}=\alpha\kappa=\left[f^{\prime}(0)\right]^{-\frac{1}{D-2}}\kappa$. After the conformal transformation, the null affine parameter $v^{E}$ on the horizon in the Einstein frame is related to that in the original frame $v$ by[12]
$\frac{\mathrm{d}v^{E}}{\mathrm{d}v}=c\left[f_{R}(x,v)\right]^{\frac{2}{D-2}},$ (42)
where $c$ is an arbitrary constant and $f_{R}(x,v)$ is the value of $f^{\prime}(\varphi)=f^{\prime}(R)$ at $(x,v)$ on the horizon. As a result,
$v^{E}=c\int_{0}^{v}\mathrm{d}v^{\prime}\left[f_{R}(x,v^{\prime})\right]^{\frac{2}{D-2}}.$ (43)
Since $0=\partial_{\tau}f^{\prime}(R)=\kappa v\partial_{v}f_{R}(x,v)$ in the stationary background, $f_{R}(x,v)$ does not depend on $v$, so $v^{E}$ is proportional to $v$ on the horizon of the stationary black hole.
IV.2 The physical process first law
Since null geodesics are conformally invariant and the Einstein equation holds for $g^{E}_{ab}$ in the Einstein frame, the derivation of Sec. III is still tenable for $f(R)$ gravity in the Einstein frame. Let
@@ IV.2 The physical process first law
denote the sum of the stress-energy tensor of the external matter fields in the Einstein frame $T^{E}_{ab}=-\frac{2}{\sqrt{-g^{E}}}\frac{\delta S_{M}}{\delta g_{E}^{ab}}=T_{ab}/f_{R}(x,v)$ as well as that of the auxiliary scalar field. Notice that
@@
where we have used the fact that $k_{E}^{a}\partial_{a}\phi=0$ and $\left.g^{E}_{ab}k^{a}k^{b}\right|_{\mathcal{H}^{+}}=0$ in the stationary background as well as the gauge condition $\delta g_{ab}k_{E}^{a}k_{E}^{b}\overset{\mathcal{H}^{+}}{=}0$ for the perturbations. Thus, only the variation of the external matter fields contributes to the first-order variation of $\mathcal{T}_{ab}$. For perturbations satisfying the gauge conditions listed in Sec. II that fix $k_{E}^{a}$ on the horizon, the non-stationary physical process first law for $f(R)$ gravity in the Einstein frame reads
@@
Next we would like to recast the physical process first law and the formula for the dynamical black hole entropy in the original frame. As the perturbation performed on the black hole is non-stationary, the gauge conditions that fix $k_{E}^{a}$ will not generally fix $k^{a}$ in the original frame. Since for non-stationary perturbations, $f_{R}(x,v)$ is not necessary a constant along the null geodesics on $\mathcal{H}^{+}$ and $v^{E}$ presumably would not be proportional to $v$ accordingly. While as mentioned in Sec. II, the essence of those gauge conditions lies in how to identify spacetime points in two slightly different spacetimes. Those gauge conditions that fix $k_{E}^{a}$ under the perturbations imply that we correspond points with the same $v^{E}$ in the two spacetimes. To rewrite (46) and (47) in the original frame, in what follows we are going to compare spacetime points sharing the same $v$ on the horizon, and replace $v^{E}$ with $v$ in the formula of $S^{E}_{\text{dyn}}$.
@@
Without loss of generality, in the following discussion we set $v^{E}=v$ in the stationary background, and perform the conformal transformation (37) on the perturbed geometry $g_{ab}+\delta g_{ab}$ to obtain the non-stationary black hole solution in the Einstein frame. Then $v^{E}=v+\delta V(x,v)$ on the horizon according to (43) in the perturbed geometry, where $\delta V(x,v)$ is a correction function derived from expanding the conformal factor $\omega^{2}=\left[f_{R}(x,v)\right]^{\frac{2}{D-2}}$ with respect to the non-stationary perturbation of the metric, and it is of first order in perturbation theory. To begin with, let $F(x,v)$ represent a quantity on the horizon, such as the entropy $S_{\text{dyn}}$ or the component of the stress-energy tensor $T_{vv}(x,v)$, which is stationary in the background, and let $\tilde{F}(x,v)$ denote the corresponding quantity in the perturbed spacetime, then $\delta F(x,v)=\tilde{F}(x,v)-F(x,v)$ does not change to first order as we switch from $v^{E}-$identification to $v-$identification, because
$\delta_{v^{E}}F-\delta_{v}F=\tilde{F}(x,v-\delta V)-\tilde{F}(x,v)\sim-\partial_{v}\tilde{F}\ \delta V=\mathcal{O}(\epsilon^{2}),$ (48)
@@
where $\delta_{v}$ stands for the comparison between points with the same $v$ in the stationary background and in the perturbed spacetime, and the same connotation holds for $\delta_{v^{E}}$. We have also used that $\partial_{v}\tilde{F}$ is of first order in the perturbed spacetime as $F$ is stationary in the unperturbed background. Therefore (46) does not change to first order under the reidentification of spacetime points. In addition, although the difference $\delta V$ between $v$ and $v^{E}$ leads to first-order corrections to the integral over $v$ and to the null generator $k^{a}$ in the perturbed geometry, the right-hand side of (46) will not change to first order if we replace $k^{a}$ by $k_{E}^{a}$ and replace $\int_{v^{E}_{1}}^{v^{E}_{2}}\mathrm{d}v^{E}$ by $\int_{v_{1}}^{v_{2}}\mathrm{d}v$, since $\delta T^{E}_{ab}$ is a first-order small quantity and this substitution leads to second-order corrections. Moreover, in the perturbed geometry,
@@
$\begin{split}S^{E}_{\text{dyn}}[\mathcal{C}(v^{E})]&=\frac{1}{4G_{D}}\left(1-v^{E}\frac{\mathrm{d}}{\mathrm{d}v^{E}}\right)\int_{\mathcal{C}}\mathrm{d}^{D-2}x\sqrt{\gamma^{E}(x,v^{E})}\\ &=\frac{1}{4G_{D}}\int_{\mathcal{C}}\mathrm{d}^{D-2}x\left[1-\left(v+\delta V\right)\left(1-\frac{\partial\delta V}{\partial v^{E}}\right)\frac{\partial}{\partial v}\right]\sqrt{\gamma}f^{\prime}(R)\\ &=\frac{1}{4G_{D}}\left(1-v\frac{\mathrm{d}}{\mathrm{d}v}\right)\int_{\mathcal{C(v)}}\mathrm{d}^{D-2}x\sqrt{\gamma}f^{\prime}(R)+\mathcal{O}(\epsilon^{2}),\end{split}$ (49)
@@
Our results agree with those obtained by means of the Noether charge method[1, 9]. Moreover, if the stress-energy tensor of the external matter field satisfies the null energy condition $\delta T_{ab}k^{a}k^{b}\geq 0$, the dynamical black hole entropy will obey the linearized second law $\Delta\delta S_{\text{dyn}}\geq 0$ under the perturbation.
IV.3 The comparison first law